KTeV Shift Summary - DAY 12/02/99 (Thur.)

Earl C. Swallow (earls@elmhurst.edu)
Thu, 2 Dec 1999 15:37:50 -0600 (CST)


Shift Summary for DAY shift, 12/02/99 (Thur.)

On Shift: Theo Alexopolous and Earl Swallow + ........

Hours of beam available at all intensities: 0.0 (out of 8.5)
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

TeV down for studies, then down for D4 "disaster" (aka "failed
component").

========================================================================
Runs ended this shift:

None.

Tapes brought to FCC this shift:

UPA,B,C208-213, UPS071-072 (20 tapes).

========================================================================
Accesses during shift:

Many, while beam is off. We are now in "open access" mode, but
shift crew should continue to control access.

========================================================================
Overview of shift:

Some beam-off work.

========================================================================
Description of Shift:

07:00 DAY Shift Begins in the wee hours of the morning.
Very quiet.

07:40 Leonard comes by to see if we need emergency action
on CsI air system. Main BryAir is producing WARM dry
air and then backup system is doing its best to cool
it. They have a repair man picking up parts and
scheduled to arrive here about 9:00 AM. We are "dead
but stable" and awaiting his arrival.

08:35 MCR tells us that they had a quench in D4 (which we
knew from reading their e-log). They may have lost
a magnet or magnets. If so, this means being off 4-5
days! More information to come at about 10:00.

09:00 Julius from Beams Division shuts the NM4AN off.

09:38 RickF calls with the bad news. We will be off for
about 7 days -- UGH! We informed Run Manager.

10:30 Spill signal problems hinder DAQ startup. MCR
signals gone and our simulator is screwed up.
RickF helps get "real" signals back from MCR, and
JasonL gets ours fixed (T6 was ORed with ours).

10:40 With authorization from Run Manager, we turn off
power to audio alarm speaker (and attach warning
note). It must be turned back on when needed.

11:14 Christmas lights (well, a few) installed in KTeV
Control Room.

15:00 WOW! Is it ever QUIET out here!

15:25 BryAir repair still in progress. Relay and fuses
replaced, and bad fan motor discovered -- repairman
is hunting part. Leonard is on top of it and will
keep us informed.

KTEV3 home disk is at 91%!!!! Disk hogs need to
clean up your files!!!!!

15:30 DAY Shift ends with a dull thud.

KTeV Logbook XXIII, pages 40-41.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shifts

Deal with disk hogs on KTEV3 home disk.

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)